# Clutch Control: An Attention-based Combinatorial Bandit for Efficient Mutation in JavaScript Engine Fuzzing

### v1

@@ 1 Introduction
JavaScript is commonly used across the web, with a consistent year-on-year increase [1, 2]. This prevalence is in part due to frameworks that allow for easy and lightweight deployment such as jQuery, React, and core-js. JavaScript is a dynamically typed, prototype based, object oriented language, that allows for flexibility of types and properties during runtime as the code executes. As such, it is used for a variety of tasks in modern web sites, including animations, logging, content loading, API calls. Due to the large volume of JavaScript on the web, and its varied functionality, JavaScript engines use Just-In-Time (JIT) compilers. These JIT compilers have become a focus of security concerns, due to the rising number of vulnerabilities they contain [3].
Fuzzing is one of the most common and efficient methodologies to find vulnerabilities in software [4]. Broadly, these approaches generate new JavaScript and then mutate test cases to change the structure, syntax, or behaviour of the test case, this is in the hope of triggering new paths in the JavaScript engine and find potential vulnerabilities or bugs. Fuzzing of JavaScript engines has spawned different methodologies ranging from neural networks for code generation [5] to heuristics [3]. However, while they focus on developing new mutations, testing heuristics, or oracles, they apply mutations at random locations within test cases. This behaviour reduces the number of successful test cases, and leads to inefficient search of system under test.
@@
We postulate it is possible to learn where to perform such a mutation, instead of random selection. The problem of selecting a given location in a JavaScript program to mutate is similar to that of contextual Combinatorial Bandits (CBs). At each timestep a representation of a test case can be observed (the contexts of arms) and locations are selected (selection of a subset of arms). The test case is immediately run and feed and feedback on the performance of the test case is retrieved (reward is received based on select arms). However, to apply a CB approach to such a task is challenging. First, the CBs must generalise quickly to the large state-action space of potential test cases generated: JavaScript fundamentals are defined in over 800 pages of technical detail in the 15th ECMA standards [6]. Second, the arms of the CB are volatile: JavaScript test cases are of variable size and structure. Third, the exploration of arms must be dynamic: random selection of testcase locations should change during testing of the JavaScript engines.
@@ 2 Background
Deep Neural Networks (DNNs) have been applied to contextual bandit problems. One of the first was NeuralUCB, which used a DNN to approximate the reward in the Upper Confidence Bound (UCB) algorithm [7]. This was closely followed the use of DNNs for Thompson Sampling (TS) [8], which introduces a posterior over the neural network to approximate the reward instead of direct approximation. Contextual CBs are a special class of bandit where a bandit selects multiple arms at each round. The CB problem is closely related to the one considered in this paper. Neural CBs were proposed by Hwang et al. [9], and are applied to both UCB (CN-UCB) and TS (CN-TS) settings. Both CN-UCB and CN-TS sample directly from the reward predicted by a DNN to select the arms to play. Volatile CBs (VCBs) consider the case where the number of arms can change with the task that is being conducted. To the best of our knowledge, the DNNs haven’t been applied to volatile bandit conditions due to the variable number of arms. The first work to consider VCB problems was CC-MAB [10] an online policy that is formed by creating identical hypercubes that exist within the space of possible arms. A similar approach is taken by ACC-UCB [10], that adaptively discretises the hypercubes to estimate outcomes of different arms. As a result, ACC-UCB functions only in a two-dimensional feature space of arms.
@@
The goal of fuzzing JavaScript engines is to trigger abnormal behaviour, corresponding to a bug or vulnerability. Fuzzing tools either tend to generate new test cases or mutate existing ones, at random or following some predefined heuristics. Code coverage is a common metric to evaluate the performance of the mutated test cases, providing a practical estimate of the portion of the search space explored, such fuzzers are referred to as greybox. However, only a small fraction of test cases will find bugs in the test program, leading to a high number of executed tests [11]. Hence, a key objective for practical fuzzing solutions is to increase the efficiency of bug finding by reducing the volume of test cases that do not result in new behaviour or do not pass input validation [12].
There has been no prior work in fuzzing to learn where to perform mutations in a test case. Limpanukorn et al. [13] use a tree-like representation of a test case to deterministically perform mutations at locations dependent on shared ancestor and sibling nodes. However, this approach is developed for low-level language representations, and not for JavaScript. Orthogonal works have focused on different aspects of fuzzing. Weissberg et al. [14] examined different methods to select areas in the target to fuzz in directed fuzzing [15, 16, 17]. Zhao et al. [18] investigated the scheduling of seeds using Monte-Carlo Tree Search approach. Lemieux et al. [19] mask locations at random, then use a language model to generate a mutation. The model is then fine-tuned via reinforcement learning to improve testing. While these works investigate selection of target code, scheduling mutators, or scheduling seeds, Clutch learns to select where in the test case to perform mutations in greybox fuzzing.
More recent research has focused specifically on the JIT compiler to target the deeper functionality and harder-to-find vulnerabilities contained within. Several of these approaches are based on Fuzzilli [20], which provided modular and customisable templates for generating JavaScript for particular engines; in addition to mutations designed to trigger JIT compilation. OptFuzz [21], builds on the Fuzzilli architecture to use optimisation paths as a performance metric. The authors of OptFuzz use an approximation of the optimisation path coverage as feedback to guide seed scheduling and preservation.
@@
RL techniques have been developed to automate many different types of testing. One of the first works to investigate RL for automating fuzzing was from Böttinger et al. [22] which focused on fuzzing C compliers. This has lead to many works that use RL for fuzzing different compilers [23, 24, 25]. Zheng et al. [26] used an RL agent with an intrinsic reward function to increase coverage in web crawling. While Tsingenopoulos et al. [27], used an alternative deep RL model to bypass captchas in web applications. RL has been used for a number of different security focused testing applications, including SQL injection [28, 29], cross site scripting [30, 31], denial-of-service [32], and cache timing attacks [33].
@@
Bandits have been used for software testing for seed scheduling. Yue et al. [34] propose a variant of an adversarial Multi-Armed Bandit (MAB), which models seed selection using the estimated reward probability. Wang et al. [35] use a similar problem formulation using an MAB, they further consider feedback in the form of hierarchical feedback on coverage of test cases. More recently Luo et al. [36] use Thompson Sampling (TS) for seed selection, with a sparse reward function. McFadden et al. [37] use PPO as a one-step MDP to classify malware samples additionally combining active learning, rejection.
@@
Fuzzing has become the de facto approach to find bugs and vulnerabilities in JavaScript engines, leading to a number of works on this. LangFuzz [38], creates an Abstract Syntax Tree (AST) of JavaScript functions, selecting nodes to mutate at random. IFuzzer [39] makes use of a similar strategy, but improves effectiveness with genetic programming, and a fitness function from greybox feedback on the test case. Montage [5] uses Natural Language Processing (NLP) to generate JavaScript subtrees at random locations in a test case’s AST. DIE [3] is a mutation based fuzzer that aims to preserve type and structural properties of test cases. CodeAlchemist [40], reduces test cases to fragments, using static and dynamic analysis to infer types, and then reassemble test cases. In both Superion [41], and SkyFire [42], the authors use grammar based methods for mutation and generation, respectively. Nautilus [43] takes this a stage further: generating its own test cases and then mutating their ASTs. While Salls et al. [44] use a ‘token level’ approach to perform mutations at a higher level, to avoid using grammars.
@@
Fuzzilli [20] is a greybox fuzzer targeting crashes in the JIT compiler of JS engines. Fuzzilli leverages coverage feedback and semantic correctness preservation in order to improve its ability to fuzz. Several approaches are based on Fuzzilli, due to the modular and customisable templates for generating JavaScript for particular engines.
FuzzJIT [45] is a tool based on the Fuzzilli architecture, but develops JIT-specific oracles and triggers. Specifically, FuzzJIT wraps each generated function in a wrapper that can be called repeatedly to trigger JIT compilation. The wrapped output can then be compared before and after JIT compilation to determine errors that result form the JIT. FuzzJIT implements its oracle to perform type-specific comparisons, using ‘deep equality’ to find if there are differences in objects, arrays, or individual values, and in such cases reporting an error.
JIT-Picker [46] uses differential testing on the backbone of Fuzzilli to detect subtle non-crash bugs. It uses both the interpreter and the JIT compiler as separate bug oracles, using the fact that both demand the interpreted and compiled code to perform computations in a strict and substitutable way. This abstracts away from implementation-dependent behaviour, allowing JIT-Picker to perform differential testing in individual engines without need to instrument the JavaScript code being tested.
@@ 3.1 Challenges in targeting mutations
Mutational JavaScript engine fuzzers traditionally focus on developing new ways to perform mutations [5, 41, 38, 3], or new oracles to detect bugs and vulnerabilities [20, 46, 21]. Mutations introduce diversity in the testing set, which may increase coverage, and find bugs. In comparison, the problem of where to apply these mutations has been overlooked, and they are typically applied at random locations. This may be in part due to the difficulty of disentangling the effects on performance of the applied mutation operator and the chosen location. We detail below two key challenges in learning where to perform mutations.
@@
Mutation based fuzzers, attempt to reduce to problem of creating test cases for fuzzing JavaScript engines by applying mutations to existing valid JavaScript. Conversely some generational based fuzzers use a constrained set of rules, or grammars, to increase the proportion of valid samples. Yet, despite this, the search space of test cases that cover new paths or trigger bugs remains vanishingly small compared to the space of syntactically valid JavaScript [40]. This can be considered a challenging search problem: how to mutate the test cases to find the ‘interesting’ test cases that trigger new areas, or behaviours of the JavaScript engines.
@@
Finding the ‘best place’ (or at least a ‘good place’) for a mutation requires identifying locations, assigning them preference, and resolving conflicting candidates. The design and implementation of appropriate heuristics would require an expansive knowledge of JavaScriptand its 800 pages of technical detail as defined in 15th ECMA standards [6].
@@ 4.1 Motivation
In the traditional MAB actions (selecting an arm) and observations do not require any kind of association. This means that in stationary problems the bandit learns the single ‘best action’, and in non-stationary tasks it learns the ‘best action’ over time [47]. However, test cases have different structures and, as in the example in Listing 1, will have different optimal locations to perform a given mutation. Thus we can consider each test case as a separate task, to optimise for coverage, bugs, or testing deeper into the JavaScript engine. As we fuzz, generated test cases will be included in the corpus increasing the number of tasks the agent has to learn. Therefore, it is important that we provide a context of the current test case to the bandit, enabling the association of actions to observations. As a result we need a contextual bandit that understands the subtleties of the JavaScript test cases.
@@ 4.2.1 Model Architecture
To handle the representation of JavaScript that forms arms and contexts for Clutch we use a deep learning architecture. Specifically we use a sequence-to-sequence model with an attention mechanism in the form of a Pointer Network [48]. Using a Gated Recurrent Unit (GRU) model as an encoder-decoder allows the encoder and decoder states and , to be used to compute an attention vector over the locations available at each timestep (), where parameters , and, are learnable.
@@
The last challenge is the rate of exploration that Clutch should use to learn over time. In deep Reinforcement Learning (RL) models this often handled by hyperparameter , which indicates the rate of selecting random actions. Conversely, deep MABs sample over a Gaussian distribution based on either the network parameters, or estimated rewards [8, 9]. However, hyperparameter search and sampling network parameters is computationally expensive. While it is possible to sample over the estimated rewards directly, this may not be sufficiently optimistic, resulting in lack of exportation and poor performance.
Instead, we can place a prior over the network, that can be tuned using Concrete Dropout [49]. In doing so, dropout parameter can be treated as a parameter to be trained via gradient descent. Thus, the exploration-exploitation trade-off can be controlled via re regularisation parameters . As training increases and more data is seen, the dropout regularisation loss reduces as the uncertainty on the weights reduces. This reduces the exploration, resulting in more exploitation. Equally, the network can have higher uncertainty in the weights which will increase the regularisation, effectively increasing the exploration. For more detailed explanation of Concrete Dropout, see Appendix B. Thus, using Eq. 1, we can form the output variables of Clutch. The encoder and decoder outputs are passed through the respective hidden layers , , with Concrete Dropout applied. This produces the attention vector , which represents the estimated rewards of each location, used to select the location as in Algorithm 1. The model also estimates the log variance, , used in a heteroskedastic loss as in Eq. 3.
@@ 4.4 Reward
Prior work from Li et al. [23] studied the use of several key metrics for rewards when fuzzing a simple compiler, including: coverage, syntactic validity, and complexity of test cases. Later work from Bates et al. [50] and Foley and Maffeis [51] have discussed the difficulty in designing rewards for real-world tasks, arguing for rewards consistent across training to ensure that diminishing returns do not effect the learnt policy. This is particularly relevant when using coverage based rewards where a) the total possible coverage can be quite large, and b) the majority of coverage is often found early in testing. Thus, we design a reward, for Clutch that is consistent across training:
@@
Finally, Eq 4 rewards for coverage using a term that incentivises the agent to explore new branches. At each timestep we compute an updated running variance of each branch () representing this as a vector: [52]. In doing so we can estimate the variance of each branch without retaining all the historic branch coverage. We then take the difference in variance between time and , isolating the difference in variance (novelty in coverage) from the test case at time . The maximum value indicates the largest difference in coverage compared to the previous timestep, i.e. where branches which are less frequently covered. We then take the absolute value to reward for changing the variance. Meaning that Clutch is incentivised not to hit the same branches frequently, but instead to explore other branches, including lower frequency branches. Compared with traditional CBs where a reward is given per arm, we reward for the set of arms, as the reward is based on the overall test case performance, not individual arms.
@@
We implement Clutch in Python 3.11 using PyTorch. We use the PythonKit [53] framework to interface Clutch with the fuzzers above (implemented in Swift), so they can share JavaScript representations, coverage information, and locations. Locations in each of these models are represented by the IR instructions found in each of the three fuzzers we place Clutch into. We run experiments on two identical machines running Ubuntu 24.04 with Intel i7 8700k and 64GB RAM.

### v2

@@ 1 Introduction
JavaScript is commonly used across the web, with a consistent year-on-year increase [http_archive_et_al_web_2023, http_archive_et_al_web_2022]. This prevalence is in part due to frameworks that allow for easy and lightweight deployment such as jQuery, React, and core-js. JavaScript is a dynamically typed, prototype based, object oriented language, that allows for flexibility of types and properties during runtime as the code executes. As such, it is used for a variety of tasks in modern web sites, including animations, logging, content loading, API calls. Due to the large volume of JavaScript on the web, and its varied functionality, JavaScript engines use Just-In-Time (JIT) compilers. These JIT compilers have become a focus of security concerns, due to the rising number of vulnerabilities they contain [park_fuzzing_2020].
Fuzzing is one of the most common and efficient methodologies to find vulnerabilities in software [zalewski_american_2007]. Broadly, these approaches generate new JavaScript and then mutate test cases to change the structure, syntax, or behaviour of the test case, this is in the hope of triggering new paths in the JavaScript engine and find potential vulnerabilities or bugs. Fuzzing of JavaScript engines has spawned different methodologies ranging from neural networks for code generation [lee_montage_2020] to heuristics [park_fuzzing_2020]. However, while they focus on developing new mutations, testing heuristics, or oracles, they apply mutations at random locations within test cases. This behaviour reduces the number of successful test cases, and leads to inefficient search of system under test.
@@
[…]
@@ 2 Background
Deep Neural Networks (DNNs) have been applied to contextual bandit problems. One of the first was NeuralUCB, which used a DNN to approximate the reward in the Upper Confidence Bound (UCB) algorithm [zhou_neural_2020]. This was closely followed the use of DNNs for Thompson Sampling (TS) [zhang_neural_2021], which introduces a posterior over the neural network to approximate the reward instead of direct approximation. Contextual CBs are a special class of bandit where a bandit selects multiple arms at each round. The CB problem is closely related to the one considered in this paper. Neural CBs were proposed by Hwang et al. [hwang_combinatorial_2023], and are applied to both UCB (CN-UCB) and TS (CN-TS) settings. Both CN-UCB and CN-TS sample directly from the reward predicted by a DNN to select the arms to play. Volatile CBs (VCBs) consider the case where the number of arms can change with the task that is being conducted. To the best of our knowledge, the DNNs haven’t been applied to volatile bandit conditions due to the variable number of arms. The first work to consider VCB problems was CC-MAB [nika_contextual_2020] an online policy that is formed by creating identical hypercubes that exist within the space of possible arms. A similar approach is taken by ACC-UCB [nika_contextual_2020], that adaptively discretises the hypercubes to estimate outcomes of different arms. As a result, ACC-UCB functions only in a two-dimensional feature space of arms.
@@
The goal of fuzzing JavaScript engines is to trigger abnormal behaviour, corresponding to a bug or vulnerability. Fuzzing tools either tend to generate new test cases or mutate existing ones, at random or following some predefined heuristics. Code coverage is a common metric to evaluate the performance of the mutated test cases, providing a practical estimate of the portion of the search space explored, such fuzzers are referred to as greybox. However, only a small fraction of test cases will find bugs in the test program, leading to a high number of executed tests [bohme_coverage-based_2016]. Hence, a key objective for practical fuzzing solutions is to increase the efficiency of bug finding by reducing the volume of test cases that do not result in new behaviour or do not pass input validation [wang_systematic_2020].
There has been no prior work in fuzzing to learn where to perform mutations in a test case. Limpanukorn et al. [limpanukorn_fuzzing_2024] use a tree-like representation of a test case to deterministically perform mutations at locations dependent on shared ancestor and sibling nodes. However, this approach is developed for low-level language representations, and not for JavaScript. Orthogonal works have focused on different aspects of fuzzing. Weissberg et al. [weissberg_sok_2024] examined different methods to select areas in the target to fuzz in directed fuzzing [lyu_mopt_2019, assiri_software_2024, srivastava_one_2022]. Zhao et al. [zhao_alphuzz_2022] investigated the scheduling of seeds using Monte-Carlo Tree Search approach. Lemieux et al. [eom_fuzzing_2024] mask locations at random, then use a language model to generate a mutation. The model is then fine-tuned via reinforcement learning to improve testing. While these works investigate selection of target code, scheduling mutators, or scheduling seeds, Clutch learns to select where in the test case to perform mutations in greybox fuzzing.
More recent research has focused specifically on the JIT compiler to target the deeper functionality and harder-to-find vulnerabilities contained within. Several of these approaches are based on Fuzzilli [gros_fuzzilli_2023], which provided modular and customisable templates for generating JavaScript for particular engines; in addition to mutations designed to trigger JIT compilation. OptFuzz [wang_optfuzz_2024], builds on the Fuzzilli architecture to use optimisation paths as a performance metric. The authors of OptFuzz use an approximation of the optimisation path coverage as feedback to guide seed scheduling and preservation.
@@
RL techniques have been developed to automate many different types of testing. One of the first works to investigate RL for automating fuzzing was from Böttinger et al. [bottinger_deep_2018] which focused on fuzzing C compliers. This has lead to many works that use RL for fuzzing different compilers [li_alphaprog_2022, li_fuzzboost_2022, sablotny_reinforcement_2023]. Zheng et al. [zheng_automatic_2021] used an RL agent with an intrinsic reward function to increase coverage in web crawling. While Tsingenopoulos et al. [tsingenopoulos_captcha_2022], used an alternative deep RL model to bypass captchas in web applications. RL has been used for a number of different security focused testing applications, including SQL injection [erdodi_simulating_2021, wahaibi_sqirl_2023], cross site scripting [lee_link_2022, foley_haxss_2022], denial-of-service [mcfadden_wendigo_2024], and cache timing attacks [luo_autocat_2023].
@@
Bandits have been used for software testing for seed scheduling. Yue et al. [yue_ecofuzz_2020] propose a variant of an adversarial Multi-Armed Bandit (MAB), which models seed selection using the estimated reward probability. Wang et al. [wang_reinforcement_2021] use a similar problem formulation using an MAB, they further consider feedback in the form of hierarchical feedback on coverage of test cases. More recently Luo et al. [luo_make_2024] use Thompson Sampling (TS) for seed selection, with a sparse reward function. McFadden et al. [mcfadden2025drmddeepreinforcementlearning] use PPO as a one-step MDP to classify malware samples additionally combining active learning, rejection.
@@
Fuzzing has become the de facto approach to find bugs and vulnerabilities in JavaScript engines, leading to a number of works on this. LangFuzz [christian_fuzzing_2012], creates an Abstract Syntax Tree (AST) of JavaScript functions, selecting nodes to mutate at random. IFuzzer [veggalam_ifuzzer_2016] makes use of a similar strategy, but improves effectiveness with genetic programming, and a fitness function from greybox feedback on the test case. Montage [lee_montage_2020] uses Natural Language Processing (NLP) to generate JavaScript subtrees at random locations in a test case’s AST. DIE [park_fuzzing_2020] is a mutation based fuzzer that aims to preserve type and structural properties of test cases. CodeAlchemist [han_codealchemist_2019], reduces test cases to fragments, using static and dynamic analysis to infer types, and then reassemble test cases. In both Superion [wang_superion_2019], and SkyFire [wang_skyfire_2017], the authors use grammar based methods for mutation and generation, respectively. Nautilus [aschermann_nautilus_2019] takes this a stage further: generating its own test cases and then mutating their ASTs. While Salls et al. [salls_token-level_2021] use a ‘token level’ approach to perform mutations at a higher level, to avoid using grammars.
@@
Fuzzilli [gros_fuzzilli_2023] is a greybox fuzzer targeting crashes in the JIT compiler of JS engines. Fuzzilli leverages coverage feedback and semantic correctness preservation in order to improve its ability to fuzz. Several approaches are based on Fuzzilli, due to the modular and customisable templates for generating JavaScript for particular engines.
FuzzJIT [wang_fuzzjit_nodate] is a tool based on the Fuzzilli architecture, but develops JIT-specific oracles and triggers. Specifically, FuzzJIT wraps each generated function in a wrapper that can be called repeatedly to trigger JIT compilation. The wrapped output can then be compared before and after JIT compilation to determine errors that result form the JIT. FuzzJIT implements its oracle to perform type-specific comparisons, using ‘deep equality’ to find if there are differences in objects, arrays, or individual values, and in such cases reporting an error.
JIT-Picker [bernhard_jit-picking_2022] uses differential testing on the backbone of Fuzzilli to detect subtle non-crash bugs. It uses both the interpreter and the JIT compiler as separate bug oracles, using the fact that both demand the interpreted and compiled code to perform computations in a strict and substitutable way. This abstracts away from implementation-dependent behaviour, allowing JIT-Picker to perform differential testing in individual engines without need to instrument the JavaScript code being tested.
@@ 3.1 Challenges in targeting mutations
Mutational JavaScript engine fuzzers traditionally focus on developing new ways to perform mutations [lee_montage_2020, wang_superion_2019, christian_fuzzing_2012, park_fuzzing_2020], or new oracles to detect bugs and vulnerabilities [gros_fuzzilli_2023, bernhard_jit-picking_2022, wang_optfuzz_2024]. Mutations introduce diversity in the testing set, which may increase coverage, and find bugs. In comparison, the problem of where to apply these mutations has been overlooked, and they are typically applied at random locations. This may be in part due to the difficulty of disentangling the effects on performance of the applied mutation operator and the chosen location. We detail below two key challenges in learning where to perform mutations.
@@
Mutation based fuzzers, attempt to reduce to problem of creating test cases for fuzzing JavaScript engines by applying mutations to existing valid JavaScript. Conversely some generational based fuzzers use a constrained set of rules, or grammars, to increase the proportion of valid samples. Yet, despite this, the search space of test cases that cover new paths or trigger bugs remains vanishingly small compared to the space of syntactically valid JavaScript [han_codealchemist_2019]. This can be considered a challenging search problem: how to mutate the test cases to find the ‘interesting’ test cases that trigger new areas, or behaviours of the JavaScript engines.
@@
[…]
@@ 4.1 Motivation
In the traditional MAB actions (selecting an arm) and observations do not require any kind of association. This means that in stationary problems the bandit learns the single ‘best action’, and in non-stationary tasks it learns the ‘best action’ over time [sutton_reinforcement_2018]. However, test cases have different structures and, as in the example in Listing 1, will have different optimal locations to perform a given mutation. Thus we can consider each test case as a separate task, to optimise for coverage, bugs, or testing deeper into the JavaScript engine. As we fuzz, generated test cases will be included in the corpus increasing the number of tasks the agent has to learn. Therefore, it is important that we provide a context of the current test case to the bandit, enabling the association of actions to observations. As a result we need a contextual bandit that understands the subtleties of the JavaScript test cases.
@@ 4.2.1 Model Architecture
To handle the representation of JavaScript that forms arms and contexts for Clutch we use a deep learning architecture. Specifically we use a sequence-to-sequence model with an attention mechanism in the form of a Pointer Network [vinyals_pointer_2017]. Using a Gated Recurrent Unit (GRU) model as an encoder-decoder allows the encoder and decoder states and , to be used to compute an attention vector over the locations available at each timestep (), where parameters , and, are learnable.
@@
The last challenge is the rate of exploration that Clutch should use to learn over time. In deep Reinforcement Learning (RL) models this often handled by hyperparameter , which indicates the rate of selecting random actions. Conversely, deep MABs sample over a Gaussian distribution based on either the network parameters, or estimated rewards [zhang_neural_2021, hwang_combinatorial_2023]. However, hyperparameter search and sampling network parameters is computationally expensive. While it is possible to sample over the estimated rewards directly, this may not be sufficiently optimistic, resulting in lack of exportation and poor performance.
Instead, we can place a prior over the network, that can be tuned using Concrete Dropout [gal_concrete_2017]. In doing so, dropout parameter can be treated as a parameter to be trained via gradient descent. Thus, the exploration-exploitation trade-off can be controlled via re regularisation parameters . As training increases and more data is seen, the dropout regularisation loss reduces as the uncertainty on the weights reduces. This reduces the exploration, resulting in more exploitation. Equally, the network can have higher uncertainty in the weights which will increase the regularisation, effectively increasing the exploration. For more detailed explanation of Concrete Dropout, see Appendix B. Thus, using Eq. 1, we can form the output variables of Clutch. The encoder and decoder outputs are passed through the respective hidden layers , , with Concrete Dropout applied. This produces the attention vector , which represents the estimated rewards of each location, used to select the location as in Algorithm 1. The model also estimates the log variance, , used in a heteroskedastic loss as in Eq. 3.
@@ 4.4 Reward
Prior work from Li et al. [li_alphaprog_2022] studied the use of several key metrics for rewards when fuzzing a simple compiler, including: coverage, syntactic validity, and complexity of test cases. Later work from Bates et al. [bates_reward_2023] and Foley and Maffeis [foley_apirl_2025] have discussed the difficulty in designing rewards for real-world tasks, arguing for rewards consistent across training to ensure that diminishing returns do not effect the learnt policy. This is particularly relevant when using coverage based rewards where a) the total possible coverage can be quite large, and b) the majority of coverage is often found early in testing. Thus, we design a reward, for Clutch that is consistent across training:
@@
Finally, Eq 4 rewards for coverage using a term that incentivises the agent to explore new branches. At each timestep we compute an updated running variance of each branch () representing this as a vector: [caussinus_updating_1982]. In doing so we can estimate the variance of each branch without retaining all the historic branch coverage. We then take the difference in variance between time and , isolating the difference in variance (novelty in coverage) from the test case at time . The maximum value indicates the largest difference in coverage compared to the previous timestep, i.e. where branches which are less frequently covered. We then take the absolute value to reward for changing the variance. Meaning that Clutch is incentivised not to hit the same branches frequently, but instead to explore other branches, including lower frequency branches. Compared with traditional CBs where a reward is given per arm, we reward for the set of arms, as the reward is based on the overall test case performance, not individual arms.
@@
We implement Clutch in Python 3.11 using PyTorch. We use the PythonKit [vieito_pvieitopythonkit_2024] framework to interface Clutch with the fuzzers above (implemented in Swift), so they can share JavaScript representations, coverage information, and locations. Locations in each of these models are represented by the IR instructions found in each of the three fuzzers we place Clutch into. We run experiments on two identical machines running Ubuntu 24.04 with Intel i7 8700k and 64GB RAM.